There exists no line with 1/√3,1/√2, 1/√6 as direction cosines.
True or false with full explanation
1
Expert's answer
2020-12-15T19:43:08-0500

Solution: Given that "There exists no line with "\frac{1}{\sqrt{3}}, \frac{1}{\sqrt{2}},\frac{1}{\sqrt{6}}" as direction cosines."

Given statement is false.

Since we have "The direction cosines are related with "l^2+m^2+n^2=1,"

where "l,m,n" are direction cosines. "

Here consider "l=\frac{1}{\sqrt{3}},m=\frac{1}{\sqrt{2}} ,n=\frac{1}{\sqrt{6}}"

Consider,

"L.H.S= l^2+m^2+n^2"


"= (\frac{1}{\sqrt{3}})^2+(\frac{1}{\sqrt{2}})^2+(\frac{1}{\sqrt{6}})^2"


"= (\frac{1}{3})+(\frac{1}{2})+(\frac{1}{6})"

"= (\frac{1}{3}+\frac{1}{2})+\frac{1}{6}"

"= (\frac{2+3}{(3)(2)})+\frac{1}{6}"


"= \frac{5}{6}+\frac{1}{6}"

"= \frac{5+1}{6}"

"= \frac{6}{6}"

"= 1"


"=L.H.S"


Therefore , there exists a line with "\frac{1}{\sqrt{3}}, \frac{1}{\sqrt{2}},\frac{1}{\sqrt{6}}" as direction cosines.

This is proved that, the given statement is false.